# Predicting Case Suffixes With Activity Start and End Times: A Sweep-Line Based Approach

### v1

@@ 5 Conclusion
This paper presents an approach to predict the sequence of remaining activities in ongoing cases of a process, along with the inter-start and processing times. Our proposed approach introduce the concept of predicting suffixes of all ongoing cases in a process, in lockstep, using a sweep line-based approach. This method enables the calculation of inter-case features at each step of the prediction process. These features help in predicting waiting times, as waiting times depend on the overall workload across all ongoing cases. Furthermore, our approach predicts case suffixes using three separate models: one for the next activity, another for inter-start time, and a third for processing time. Our approach enables optimizing each model separately and predicting the end time of ongoing activity instances, given that their activity type and start times.
@@
Our evaluation show that this approach (namely MM) has better accuracy relative to existing baselines that predict the activity and timestamps via one single model (SM approaches). In particular, we observed an improvement in the control-flow metrics, attributable to the fact that we exploited the modularity of the proposed approach to incorporate an optimized method for next-activity sampling. We also observed improvements in inter-start time prediction attributable to the use of inter-case features in the proposed sweep-line method.
In future work, we plan to explore alternative architectures such as Transformers and encoder-decoder models with attention for next activity, inter-start time, and processing time prediction. We also aim to reduce reliance on teacher forcing by mixing predicted and ground-truth activities during training. Additionally, we intend to apply our approach to use cases like capacity planning and scheduling. To further assess the role of inter-case features, we plan to compare our method against state-of-the-art models that do not use such features.
@@
https://shorturl.at/RXj4d, https://shorturl.at/kJA01

### v2

@@ 5 Conclusion
This paper presents a method to predict the sequence of remaining activity instances for each ongoing case of a business process, along with the inter-start time and processing time of each activity instance. The proposed method predicts the suffixes of all ongoing cases in lockstep, using a sweep line-based approach. At each cut-off point, starting from the time of the most recently started activity instance, we estimate the label of the next activity instance in the same case, and the (inter-)start time and processing time of this activity instance. After each such next-activity prediction step, we recalculate a set of features related to resource contention and resource availability. Resource contention and availability are two key sources of waiting time in business processes. Accordingly, the proposed method builds on the hypothesis that the inclusion of such features is likely to enhance the accuracy of inter-start time predictions.
The proposed method predicts case suffixes using three separate models: one for the next activity prediction, another for inter-start time prediction, and a third for processing time prediction. This multi-model approach enables us to optimize each model separately.
@@
The experimental evaluation shows that the proposed multi-model ($M$) method achieves higher accuracy relative to existing baselines, which predict the activity and timestamps via a single model (SM approaches) and which do not incorporate resource contention and resource availability features. In particular, we observed an improvement in the control-flow accuracy metrics, attributable to the fact that we exploited the modularity of the proposed approach to incorporate an optimized method for next-activity sampling. We also observed improvements in inter-start time prediction attributable to the use of resource contention features in the proposed sweep-line method. Furthermore, we observed that when the resource availability features are integrated within the $M$ approach, this information significantly improved the performance of $M$’s variant $MM_{avail}$ for case suffix prediction. This improvement was particularly evident with respect to control flow, inter-start time, and processing time prediction. In comparison to the previous $M$ approach, $MM_{avail}$ demonstrated improved performance, highlighting the effectiveness of incorporating resource availability features into the predictive model for predicting case suffix of an ongoing case in a business process.
6 Future Work
In this paper, we developed predictive models that leverage features capturing two primary sources of waiting times in business processes: resource contention and resource availability. While these factors account for a significant portion of waiting times in business processes, previous studies by Ali et al. [3] and Lashkevich et al. [19] have shown that waiting times also arise from other sources, including batching effects, prioritization policies, and external factors beyond the scope of the process.
A promising direction for future work is to extend the proposed approach by incorporating features that explicitly model these additional sources of waiting time. For example, future approaches could capture batching behaviour by representing batch formation rules and batch sizes, account for prioritization by modelling case-level or activity-level priority schemes, and include external contextual information (e.g., calendar effects or unexpected disruptions) to better explain externally induced delays. Integrating these factors is expected to further improve the accuracy and robustness of waiting time predictions.
In addition, we plan to apply the proposed approach to practical capacity planning and scheduling use cases. In particular, the learned models could be used by process managers to evaluate and optimize alternative resource availability schedules given the current state of a running process. For instance, in scenarios where a large number of cases are pending and additional resources (e.g., workers) can be temporarily reassigned, the proposed approach could be used to answer “what-if” questions such as: Given the current workload and the availability of ten additional workers, what resource schedules would best ensure that ongoing cases meet their deadlines? Addressing such decision-support scenarios represents an important step toward operationalizing predictive process monitoring models in real-world settings.
@@ 6 Future Work
https://shorturl.at/wcu1P, https://shorturl.at/PM3aV